OpenStudy (anonymous):

12!=__________________ a) 12x1 b) 12x11x10x9! c) 12 d) 12x11x10x9x8x7x6x5x4x3x2x1 e) 12!x11! (can have more than one asnwer, it has been a while since I've done this)

OpenStudy (angela210793):

b and d

OpenStudy (anonymous):

b and d

OpenStudy (anonymous):

b

OpenStudy (anonymous):

you have to know that n! = n*(n-1)*(n-2*)(n-3).....*(n-x)! can also be written as n! = n*(n-1)! or n! = n(*n-1)*(n-2)! and so on so 12! = 12*11! 12! = 12*11*10! 12!=12*11*10*9! 12!=12*11*10*9*8*.........*2*1 therefore the answer is b and d